OpenStudy (anonymous):

Turn 12x+6y=18 into slope-intercept for.

OpenStudy (anonymous):

*form*

OpenStudy (anonymous):

is it 2x+-3

OpenStudy (anonymous):

12x + 6y = 18 Isolate y on one side. So subtract 12x from both sides. 12x + 6y = 18 -12x -12x 6y = -12x + 18 Divide each side by 6. 6y/6 = -12x/6 + 18/6 y = -2x + 3 <----- Answer. (:

OpenStudy (anonymous):

Thanks again!!!!!

OpenStudy (anonymous):

You were close!

OpenStudy (anonymous):

You're welcome. (:

OpenStudy (anonymous):

The negitive numbers like always throught me off:P

OpenStudy (anonymous):

Haha, it happens a lot to everyone! Just be careful about them.

OpenStudy (anonymous):

okay I will!!!:D